Due to Covid 19, managers of mining companies are deciding to have fewer workers, the workers size in a company has decreased over the past months. According to Chamber of Mines report, the mean workers size was 3.18 in 2019. A researcher wanted to check if the current mean workers size is less than 3.18. A sample of 900 workers taken this year by this researcher produced a mean workers size of 3.16 with a standard deviation of 0.70. Using the 0.025 significance level, can you conclude that the mean workers size has decrease since 2019?


1
Expert's answer
2020-07-06T20:14:49-0400

"H_0: \mu\geq3.18\\
H_1:\mu <3.18\\
 \alpha =0.025\implies Z_{0.025}=1.96,\\
Z=\frac{\bar x-\mu_0}{\frac{s}{\sqrt n}}\\
Z=\frac{3.16-3.18}{\frac{0.7}{\sqrt {900}}}=-0.857\\
\text{the rejection region} :\; Z<-1.96\\
\text{The decision is to accept}:\; H_0\\
\text {This means that you cannot conclude that }\\
\text{ the mean workers size has decreased since 2019 }\\"